OpenStudy (anonymous):

Harold left town driving east at 55 mph. Carol left 2 hours later driving the same route at 65 mph. How many hours after Harold left will it take Carol to catch up to Harold?

hero (hero):

d = rt Harold's Distance = 55*t Carol's Distance = 65(t-2) to get the time when their distance will be equal, set their distances equal to each other: 55t = 65(t-2) 55t = 65t-130 130 = 10t t = 13 So they will run into each other 13 hours later